Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 334     Entry time: 07 Sep 2023, 07:24
Author: Justine Schlappa, Guiseppe Mercurio 
Group: SCS 
Subject: Shift summary, September 7 after nightshift 

Achievements:
- Optimized hRIXS settings for OK edge at 145 deg scattering angle, resolution of about 50 meV
- Transferred the user sample
- Prepared for delay scans on fluorescence signal from YBCO

Problems:
- There was no internet connection at XFEL and DESY for some time
- hRIXS camera goes frequently into error
- XRD motors (STY) go frequently into error

Next steps:
- Find time zero on fluoresce signal from YBCO, test the same for the samples
- Test PP laser damage threshold

ELOG V3.1.4-7c3fd00